A motorboat goes downstream in a river and covers the distance between two coastal towns in 5 hours. It covers this distance upstream in 6 hours. If the speed of the stream is 2 km/h, find the speed of the boat in still water.
Answer by ikleyn(53763)

.
A motorboat goes downstream in a river and covers the distance between two coastal towns in 5 hours.
It covers this distance upstream in 6 hours. If the speed of the stream is 2 km/h, find the speed of the boat in still water.
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

Let "u" be the boat speed in still water, in km/h.

Then the boat speed downstream is (u+2) km/h, and the distance downstream is

5*(u+2) kilometers.  (The distance = rate*time, as you know).


The boat speed upstream is (u-2) km/h, and the distance upstream is 

6*(u-2) kilometers.


Since the distance downstream is the same as upstream, you have this equation

5(u+2) = 6(u-2).

Simplify and solve it for "u". You will get the answer u = 22 km/h.


Answer.  The speed of the boat in still water is 22 km/h.